OpenStudy (anonymous):

solve for x:2/x-2+7/x^2-4=5/x

OpenStudy (owlcoffee):

Can you use the equation editor or draw it out, it's not very clear.

OpenStudy (anonymous):

\[\frac{ 2 }{ x-2 }+\frac{ 7 }{ x^2-4 }=\frac{ 5 }{ x }\]

OpenStudy (owlcoffee):

okay, it's like this: \[\frac{ 2 }{ x-2 }+\frac{ 7 }{ x ^{2}-4 }=\frac{ 5 }{ x }\] It's a ratonal equation (why?) because we see there exists a equality and the variable "x" is situated in the denominator. Our goal is to simplify it as much as we can, so it becomes something more simple, for example, we can observe that in the 2nd denominador we have a square difference and that we can factor: \[\frac{ 2 }{ x-2 }+\frac{ 7 }{ (x+2)(x-2) }=\frac{ 5 }{ x }\]

OpenStudy (anonymous):

i only have 4/3 as an option that a fraction

OpenStudy (anonymous):

@Owlcoffee

OpenStudy (anonymous):